# The IACOB project

### v1

@@ Abstract
Aims. We aim to create an updated Hertzsprung-Russell (HR) diagram that includes a volume-limited and statistically significant sample of massive stars with luminosities and effective temperatures 14 kK. Our goal is to use this sample to investigate the extension and end of the MS, also incorporating information about projected rotational velocities () and spectroscopic binarity status.
Methods. We combine spectroscopic parameters derived with the FASTWIND stellar atmosphere code and Gaia distances to obtain estimates of stellar parameters for a sample of 876 Galactic luminous O- and B-type stars gathered within the IACOB project. We use the ALS III catalog of Galactic OB stars to select the best-suited volume-limited sample for our study. We choose the iacob-broad tool to derive estimates and review multi-epoch spectra to identify single- and double-line spectroscopic binaries (SB1, SB2).
Results. We present an HR diagram for a sample of 670 stars located within 2500 pc that has the best balance between completeness and number. We evaluate the extension of the MS in terms of the drop in the relative number of stars as a function of the effective temperature for different luminosity ranges. We find a consistent cool boundary at 22.5 kK within the full range of luminosities that we use to delineate the terminal-age main sequence (TAMS). We obtain a smooth decrease of the highest observed with along the MS band, likely limited by the critical velocity. We consider this effect combined with a lower expected fraction of stars beyond the MS as the best explanation for the lack of fast-rotating objects in the post-MS region. Our results favor low to mild initial rotation ( 150 km s-1) for the full sample along with a binary past for the well-known tail of fast-rotating stars in the distribution. The prominence of SB1 and SB2 systems within the MS band, and the 25% decrease in the relative fraction of SB1 systems when crossing the TAMS can further delineate its location.
@@
Massive stars (M 8 M⊙) represent the minority of stars in the local Universe (Salpeter 1955; Chabrier 2005). Despite their low relative number, their influence on the interstellar medium and the chemical evolution of their host galaxies is enormous (e.g., Maeder & Meynet 2000; Langer 2012; Nomoto et al. 2013; Smith 2014). Furthermore, they are among the most luminous stars in the Universe, becoming useful tools as distance indicators (Kudritzki et al. 1999, 2003), and the main stellar contributors to the brightness of high-redshift galaxies (e.g., Dijkstra & Wyithe 2007). Despite their influence and importance for many branches of astrophysics, the physical properties and processes that govern their evolution are not yet fully understood.
@@
Traditionally, blue supergiants (BSGs; B-type stars with luminosity classes I and II) were considered to be post-MS objects following MS O-type stars. However, the discovery of an overdensity of BSGs by Fitzpatrick & Garmany (1990) within the Hertzsprung gap raised questions about the true extension of the main sequence of massive stars and the true nature of BSGs. One possibility to solve the still persistent uncertainty is that a significant fraction of BSGs could be MS objects. This would imply a reconsideration of the MS extension in the evolutionary models. In this regard, the length of the MS depends not only on the stellar mass and the initial rotation rates of the stars (Meynet & Maeder 2005; Heger et al. 2005; Georgy et al. 2013), but also on different proposed internal mixing processes that can provide additional hydrogen to the star’s core (Maeder 1981, 2009; Schootemeijer et al. 2019; Martinet et al. 2021), thus extending the MS. Furthermore, massive stars are affected by strong stellar winds, which can also affect their lifetimes through the effect of mass loss (Kudritzki & Puls 2000; Vink et al. 2010; Smith 2014). Another possibility is that additional evolutionary channels contribute to the observed overdensity of BSGs. These may include stars evolving blueward, such as those that undergo a blue loop after the red supergiant (RSg) phase (Stothers & Chin 1975; Ekström et al. 2012), but also stars that have evolved through binary interactions (Sana et al. 2012; de Mink et al. 2013; Menon et al. 2024). In this entangled context, observational constraints are the key to improving the situation.
@@
Since the work by Blaha & Humphreys (1989) and Fitzpatrick & Garmany (1990) using photometric data of stars in the Galaxy and the Large Magellanic Cloud, only the study conducted by Castro et al. (2014) using a statistically significant sample of massive stars with derived spectroscopic parameters from the literature has attempted to better constrain the width of the MS. Although their sample comprised 400 Galactic OB stars, and their results evidenced the extension of the MS, the presence of observational biases and the lack of homogeneity in their analysis prevented them from quantitatively providing the location of the terminal-age main sequence (TAMS) for masses greater than 20 M⊙. Consequently, high-quality statistically significant samples of BSGs homogeneously analyzed and for which additional spectroscopic information is available (including spin rates, binary status, or surface abundances) are key to improving the current situation, which is the main goal of this work. Following Martinet et al. (2021), we define the TAMS by the position in the MS band where the evolutionary tracks reach their lowest effective temperature before momentarily returning to a hotter region (the so-called “hook”), at core hydrogen exhaustion (see also Castro et al. 2014).
@@
In this paper, we benefit from the spectroscopic observations gathered within the IACOB project (Simón-Díaz et al. 2020) and from the astrometric distances delivered by the Gaia mission (Gaia Collaboration et al. 2016, 2022; Babusiaux et al. 2023) to investigate three empirical aspects that can shed light on the extension of the MS and the location of the TAMS in the 12 – 40 M⊙ range of initial masses. These are the distribution of stars in the HR diagram in terms of density and relative number, the spin-rate properties, and the binary status.
@@
This paper is organized as follows. Section 2 presents the spectroscopic sample of stars and data used to carry out this work. Section 3 describes the methodology used to evaluate the completeness of our sample. In Sect. 4 we present the HR diagram of the stars, their empirical properties evaluated in this work, our proposed TAMS compared to the most-used evolutionary models, and a comparison with the results of Castro et al. (2014). In Sect. 5 we discuss the location of the TAMS in the HR diagram and the rotational properties of the sample, both connected with further implications for massive star evolution. Section 6 presents the main conclusions and the follow-up work.
@@ 2.1 Spectroscopic sample
The spectroscopic sample of this work comprises 876 O3- to B6-type stars that cover the ranges of 49 – 14 kK in effective temperature (), 4.2 – 1.8 dex in surface gravity (), and 6.0 – 4.3 dex in luminosity (). Of them, 678 were selected from the sample of late O- and early B-type stars presented in de Burgos et al. (2023) as stars putatively descended from stellar objects born as O-type stars (in the context of single-star evolution). This sample was later enlarged with additional observations of stars of similar types to increase the completeness (see Sect. 3). The remaining 198 objects were selected from Holgado et al. (2018, 2020, 2022). In total, 45% of the sample corresponds to stars classified as B-type supergiants and bright giants, and 35% to O-type stars. The remaining 20% corresponds to B-type stars classified as giants to dwarfs.
All stars in the sample have magnitude , and their optical spectra were collected from the IACOB spectroscopic database and the ESO public archive. In particular, they were acquired using either FIES (Telting et al. 2014, 2.5 m NOT), HERMES (Raskin et al. 2011, 1.2 m Mercator), or FEROS (Kaufer et al. 1997, 2.2 m MPG/ESO) high resolution (R = ) optical (3800 – 7000 Å) echelle spectrographs. The average signal-to-noise ratio is 130 (measured at 4500 Å). For more details, see de Burgos et al. (2023).
@@ 2.3 Spectroscopic parameters
We applied the same methodology as in de Burgos et al. (2024b) to derive spectroscopic parameters for those O9 – B6 stars not analyzed there that are included in the present sample. Briefly, the line-broadening analysis was performed using the iacob-broad tool (see Simón-Díaz & Herrero 2014), which allows to obtain estimates of the projected rotational velocity (). For the remaining spectroscopic parameters used here, namely and , the analysis was carried out using a grid of unclumped model atmospheres (i.e., not considering wind inhomogeneities) computed with the NLTE model atmosphere and the line synthesis code FASTWIND (Fast Analysis of STellar atmospheres with WINDs, v10.4.7, Santolaya-Rey et al. 1997; Puls et al. 2005; Rivero González et al. 2011; Puls et al. 2020). Although we refer to Sect. 3 in de Burgos et al. (2024b) for additional details on the properties of the grid, the main advantage of our analysis comes from the use of a statistical emulator of synthetic spectra, which is combined with a Markov chain Monte Carlo (MCMC) method to obtain the probability distribution function for the parameters mentioned above. For the 198 O-type stars from Holgado et al. (2018, 2020, 2022), was also obtained using the iacob-broad, while and were obtained using the iacob-gbat/FASTWIND tool (Simón-Díaz et al. 2011; Sabín-Sanjulián et al. 2014). We refer to those works for more details.
@@ 2.5 Radial velocity measurements
We obtained radial velocity (RV) estimates and peak-to-peak amplitude values (RVpp) in 605 stars for which multi-epoch spectra in the IACOB spectroscopic database are available. For most O-type stars, we adopted the values from Holgado et al. (2018). The methodology followed there consisted of a cross-correlation between the observed spectra and the best-fitting synthetic FASTWIND spectra for a given set of diagnostic lines. For most O9- and all B-type stars in the sample, our estimates were obtained following the methodology described in de Burgos et al. (2020). In this case, individual lists of several diagnostic lines are selected and optimized for each sub-spectral type. Then, a combination of line-fitting and sigma-clipping routines allowed us to obtain an average estimate for each spectrum. Alternatively, and only for stars with 150 km s-1, the estimates have been derived using a cross-correlation technique similar to Holgado et al. (2018). Peak-to-peak amplitudes were obtained simply as the difference between the largest difference of all the individual values.
@@ 3 Completeness analysis
For our completeness evaluation, we selected only those stars in the ALS III with B6 spectral types or earlier and with a ¡ 11, as quoted in the SIMBAD astronomical database (Weis & Bomans 2020). This was made to limit the catalog to stars with the same spectral types and magnitude range as in our sample (see also below). We also removed hypergiant, Be, and SB2+ stars identified as such in previous works, as they are not in our sample (see Sect. 2.1). Additionally, we removed missing stars identified as classical Be-stars in the BeSS (Neiner et al. 2011) and BeSOS (Arcos et al. 2021) databases, plus some other stars identified as such in SIMBAD. Last, we removed stars with a distance larger than 4000 pc in Bailer-Jones et al. (2021) or the ALS III catalog, and those with a poor distance determination, which mostly corresponds to very bright stars with a large parallax uncertainty (/ 0.15).
@@ 4.1 The Hertzsprung-Russell diagram: drop in density
Figure 2 includes three different HR diagrams showing our sample of O3 – B6 spectral type stars for which we account with reliable fundamental parameters. Each one considering stars up to a different distance. The central panel shows those stars located up to 2500 pc, that is, to our considered best distance in terms of completeness, and includes a total of 670 stars. The right panel extends the distance to 4000 pc and displays 852 of the 876 stars of the complete sample. The left panel is limited to stars up to 1500 pc, and includes 221 stars.
Regardless of the distance, the distribution of stars in the three panels shows an apparent lower density of objects below 23 – 21 kK. To better characterize this feature, Fig. 3 shows the number of stars against their effective temperature on a grid of histograms that separate stars in different distances and luminosity ranges as indicated in the panels (the latter ranges are also included in Fig. 2). Looking at the histograms for stars within 2500 pc (middle column), those panels evidence a consistent drop in the relative number of stars below 22 kK, being less steep only in the top panel of stars in the highest luminosity range. As will be discussed in Sect. 5 (see also Sect. 4.3), we interpret the drop in density as a consequence of stars leaving the MS and, therefore, outlining the location of the TAMS.
@@
Figure 3 shows that for stars with ¡ 5.0 dex the adopted position of the drop is consistent for stars within 1500 pc and 4000 pc. Regarding the stars with ¿ 5.0 dex up to 1500 pc, we argue that despite the higher degree of completeness, the statistical number of stars is not sufficient to show the drop. For stars with ¿ 5.35 dex within 4000 pc, it can be seen that the drop is somewhat shifted towards cooler temperatures. This could be the result of missing a non-negligible number of stars that are intrinsically fainter on the hot side of the drop, as mid-B-type stars are intrinsically brighter than their early-B and O-type companions (in other words, caused by the Malmquist bias; Malmquist 1922).
Regarding the distribution of stars excluded from our sample (see Sect. 2.1), we expect Be stars to only populate the lower panels of Fig. 3, since classic Be stars mostly comprise early-B MS objects with luminosity classes III to V, and to do it evenly with (see, e.g., Arcos et al. 2018). On the other hand, SB2+ systems represent an important fraction in the O-star domain. For example, Holgado et al. (2022) showed that 30% of their sample of O-type stars are SB2 systems (see also Barbá et al. 2017; Sana 2017), while in de Burgos et al. (2023, see Fig. 6) we found that the percentage among early- to mid-B-type supergiants drops below 10%. The lack of SB2+ systems could for instance explain the apparently lower number of objects above 30 kK shown in Fig. 2 (see also Sect. 5.2). In addition, the lower number of objects around 32 kK with ¿ 5.0 dex (more clearly shown in Fig. 3) could be explained considering that O9-type stars, which populate that parameter space, gather the largest number of detected SB2 systems. However, a proper diagnostic of this feature would also require a better understanding on how binary interaction products populate the HR diagram. We also note that the left-skewed distributions shown in Fig. 3, particularly for stars with ¡ 5.0 dex, partially result from the effect of stars reaching different maximum values along the ZAMS.

### v2

@@ Abstract
Aims. We aim to create an updated Hertzsprung-Russell (HR) diagram that includes a volume-limited and statistically significant sample of massive stars with luminosities and effective temperatures 14 kK. Our goal is to use this sample to investigate the extension and end of the MS while also incorporating information about projected rotational velocities () and spectroscopic binarity status.
Methods. We combined spectroscopic parameters derived with the FASTWIND stellar atmosphere code and Gaia distances to obtain estimates of stellar parameters for a sample of 876 Galactic luminous O- and B-type stars gathered within the IACOB project. We used the ALS III catalog of Galactic OB stars to select the best-suited volume-limited sample for our study. We chose the iacob-broad tool to derive estimates and reviewed multi-epoch spectra to identify single- and double-line spectroscopic binaries (SB1, SB2).
Results. We present an HR diagram for a sample of 670 stars located within 2500 pc that has the best balance between completeness and number. We evaluated the extension of the MS in terms of the drop in the relative number of stars as a function of the effective temperature for different luminosity ranges. We found a consistent cool boundary at 22.5 kK within the full range of luminosities that we used to delineate the terminal-age main sequence (TAMS). We obtained a smooth decrease of the highest observed with along the MS band, which is likely limited by the critical velocity. We consider this effect combined with a lower expected fraction of stars beyond the MS as the best explanation for the lack of fast-rotating objects in the post-MS region. Our results favor low to mild initial rotation ( 150 km s-1) for the full sample along with a binary past for the well-known tail of fast-rotating stars in the distribution. The prominence of SB1 and SB2 systems within the MS band and the 25% decrease in the relative fraction of SB1 systems when crossing the TAMS can further delineate its location.
@@
Massive stars (M 8 M⊙) represent the minority of stars in the local Universe (Salpeter 1955; Chabrier 2005). Despite their low relative number, their influence on the interstellar medium and the chemical evolution of their host galaxies is enormous (e.g., Maeder & Meynet 2000; Langer 2012; Nomoto et al. 2013; Smith 2014). Furthermore, they are among the most luminous stars in the Universe, making them useful tools as distance indicators (Kudritzki et al. 1999, 2003), and the main stellar contributors to the brightness of high-redshift galaxies (e.g., Dijkstra & Wyithe 2007). Despite their influence and importance for many branches of astrophysics, the physical properties and processes that govern their evolution are not yet fully understood.
@@
Traditionally, blue supergiants (BSGs; B-type stars with luminosity classes I and II) were considered to be post-MS objects following MS O-type stars. However, the discovery of an overdensity of BSGs by Fitzpatrick & Garmany (1990) within the Hertzsprung gap raised questions about the true extension of the main sequence of massive stars and the true nature of BSGs. One possibility that would solve the still persistent uncertainty is considering that a significant fraction of BSGs are MS objects. This would imply a reconsideration of the MS extension in the evolutionary models. In this regard, the length of the MS depends not only on the stellar mass and the initial rotation rates of the stars (Meynet & Maeder 2005; Heger et al. 2005; Georgy et al. 2013) but also on different proposed internal mixing processes that can provide additional hydrogen to the star’s core (Maeder 1981, 2009; Schootemeijer et al. 2019; Martinet et al. 2021), thus extending the MS. Furthermore, massive stars are affected by strong stellar winds, which can also affect their lifetimes through the effect of mass loss (Kudritzki & Puls 2000; Vink et al. 2010; Smith 2014). Another possibility is that additional evolutionary channels contribute to the observed overdensity of BSGs. These may include stars evolving blueward, such as those that undergo a blue loop after the red supergiant (RSg) phase (Stothers & Chin 1975; Ekström et al. 2012), but also stars that have evolved through binary interactions (Sana et al. 2012; de Mink et al. 2013; Menon et al. 2024). In this entangled context, observational constraints are the key to improving the situation.
@@
Since the work by Blaha & Humphreys (1989) and Fitzpatrick & Garmany (1990) using photometric data of stars in the Galaxy and the Large Magellanic Cloud, only the study conducted by Castro et al. (2014) using a statistically significant sample of massive stars with derived spectroscopic parameters from the literature has attempted to better constrain the width of the MS. Although their sample comprised 400 Galactic OB stars and their results evidenced the extension of the MS, the presence of observational biases and the lack of homogeneity in their analysis prevented them from quantitatively providing the location of the terminal-age MS (TAMS) for masses greater than 20 M⊙. Consequently, high-quality statistically significant samples of BSGs homogeneously analyzed and for which additional spectroscopic information is available (including spin rates, binary status, and surface abundances) are key to improving the current situation, which is the main goal of this work. Following Martinet et al. (2021), we define the TAMS by the position in the MS band where the evolutionary tracks reach their lowest effective temperature before momentarily returning to a hotter region (the so-called hook) at core hydrogen exhaustion (see also Castro et al. 2014).
@@
In this paper, we benefit from the spectroscopic observations gathered within the IACOB project (Simón-Díaz et al. 2020) and from the astrometric distances delivered by the Gaia mission (Gaia Collaboration et al. 2016, 2023; Babusiaux et al. 2023) to investigate three empirical aspects that can shed light on the extension of the MS and the location of the TAMS in the 12 – 40 M⊙ range of initial masses. These are the distribution of stars in the HR diagram in terms of density and relative number, the spin-rate properties, and the binary status.
@@
This paper is organized as follows. Section 2 presents the spectroscopic sample of stars and data used to carry out this work. Section 3 describes the methodology used to evaluate the completeness of our sample. In Sect. 4 we present the HR diagram of the stars, their empirical properties evaluated in this work, our proposed TAMS compared to the most-used evolutionary models, and a comparison with the results of Castro et al. (2014). In Sect. 5 we discuss the location of the TAMS in the HR diagram and the rotational properties of the sample, as both are connected with further implications for massive star evolution. Section 6 presents the main conclusions and the follow-up work.
@@ 2.1 Spectroscopic sample
The spectroscopic sample of this work comprises 876 O3- to B6-type stars that cover the ranges of 49 – 14 kK in effective temperature (), 4.2 – 1.8 dex in surface gravity (), and 6.0 – 4.3 dex in luminosity (). Of them, 678 were selected from the sample of late O- and early B-type stars presented in de Burgos et al. (2023) as stars putatively descended from stellar objects born as O-type stars (in the context of single-star evolution). This sample was later enlarged with additional observations of stars of similar types to increase its completeness (see Sect. 3). The remaining 198 objects were selected from Holgado et al. (2018, 2020, 2022). In total, 45% of the sample corresponds to stars classified as B-type supergiants and bright giants and 35% to O-type stars. The remaining 20% corresponds to B-type stars classified as giants to dwarfs.
All stars in the sample have magnitudes of , and their optical spectra were collected from the IACOB spectroscopic database and the ESO public archive. In particular, they were acquired using either FIES (Telting et al. 2014, 2.5 m NOT), HERMES (Raskin et al. 2011, 1.2 m Mercator), or FEROS (Kaufer et al. 1997, 2.2 m MPG/ESO) high resolution (R = ) optical (3800 – 7000 Å) echelle spectrographs. The average signal-to-noise ratio is 130 (measured at 4500 Å; for more details, see de Burgos et al. 2023).
@@ 2.3 Spectroscopic parameters
We applied the same methodology as in de Burgos et al. (2024b) to derive the spectroscopic parameters for the O9 – B6 stars not analyzed there that are included in the present sample. Briefly, the line-broadening analysis was performed using the iacob-broad tool (see Simón-Díaz & Herrero 2014), which allows one to obtain estimates of the projected rotational velocity (). For the remaining spectroscopic parameters used here, namely and , the analysis was carried out using a grid of unclumped model atmospheres (that is, not considering wind inhomogeneities) computed with the non-local thermodynamic equilibrium model atmosphere and the Fast Analysis of STellar atmospheres with WINDs line synthesis code (FASTWIND v10.4.7, Santolaya-Rey et al. 1997; Puls et al. 2005; Rivero González et al. 2011; Puls et al. 2020). Although we refer to Sect. 3 in de Burgos et al. (2024b) for additional details on the properties of the grid, the main advantage of our analysis comes from the use of a statistical emulator of synthetic spectra, which is combined with a Markov chain Monte Carlo (MCMC) method to obtain the probability distribution function for the parameters mentioned above. For the 198 O-type stars from Holgado et al. (2018, 2020, 2022), was also obtained using the iacob-broad, while and were obtained using the iacob-gbat/FASTWIND tool (Simón-Díaz et al. 2011; Sabín-Sanjulián et al. 2014). We refer to those works for more details.
@@ 2.5 Radial velocity measurements
We obtained radial velocity (RV) estimates and peak-to-peak amplitude values (RVpp) in 605 stars for which multi-epoch spectra in the IACOB spectroscopic database are available. For most O-type stars, we adopted the values from Holgado et al. (2018). The methodology followed there consisted of a cross-correlation between the observed spectra and the best-fitting synthetic FASTWIND spectra for a given set of diagnostic lines. For most O9- and all B-type stars in the sample, our estimates were obtained following the methodology described in de Burgos et al. (2020). In this case, individual lists of several diagnostic lines are selected and optimized for each sub-spectral type. Then, a combination of line-fitting and sigma-clipping routines allowed us to obtain an average estimate for each spectrum. Alternatively, and only for stars with 150 km s-1, the estimates were derived using a cross-correlation technique similar to Holgado et al. (2018). Peak-to-peak amplitudes were obtained simply as the difference between the largest difference of all the individual values.
@@ 3 Completeness analysis
For our completeness evaluation, we selected only those stars in the ALS III with B6 spectral types or earlier and with a ¡ 11, as quoted in the SIMBAD astronomical database (Weis & Bomans 2020). This was made to limit the catalog to stars with the same spectral types and magnitude range as in our sample (see also below). We also removed hypergiant, Be, and SB2+ stars identified as such in previous works, as they are not in our sample (see Sect. 2.1). Additionally, we removed missing stars identified as classical Be-stars in the BeSS (Neiner et al. 2011) and BeSOS (Arcos et al. 2021) databases, plus some other stars identified as such in SIMBAD. Last, we removed stars with a distance larger than 4000 pc in Bailer-Jones et al. (2021) or the ALS III catalog and those with a poor distance determination, which mostly corresponds to very bright stars with a large parallax uncertainty (/ 0.15).
@@ 4.1 The Hertzsprung-Russell diagram: Drop in density
Figure 2 includes three different HR diagrams showing our sample of O3 – B6 spectral type stars for which we account with reliable fundamental parameters. Each one considers stars up to a different distance. The central panel shows stars located up to 2500 pc, that is, to our considered best distance in terms of completeness, and includes a total of 670 stars. The right panel extends the distance to 4000 pc and displays 852 of the 876 stars of the complete sample. The left panel is limited to stars up to 1500 pc and includes 221 stars.
Regardless of the distance, the distribution of the stars in the three panels shows an apparent lower density of objects below 23 – 21 kK. To better characterize this feature, Fig. 3 shows the number of stars against their effective temperature on a grid of histograms that separate stars in different distances and luminosity ranges as indicated in the panels (the latter ranges are also included in Fig. 2). Looking at the histograms for stars within 2500 pc (middle column), those panels evidence a consistent drop in the relative number of stars below 22 kK, being less steep only in the top panel of stars in the highest luminosity range. As is discussed in Sect. 5 (see also Sect. 4.3), we interpret the drop in density as a consequence of stars leaving the MS and therefore outlining the location of the TAMS.
@@
Figure 3 shows that for stars with ¡ 5.0 dex the adopted position of the drop is consistent for stars within 1500 pc and 4000 pc. Regarding the stars with ¿ 5.0 dex up to 1500 pc, we argue that despite the higher degree of completeness, the statistical number of stars is not sufficient to show the drop. For stars with ¿ 5.35 dex within 4000 pc, it can be seen that the drop is somewhat shifted toward cooler temperatures. This could be the result of missing a non-negligible number of stars that are intrinsically fainter on the hot side of the drop, as mid-B-type stars are intrinsically brighter than their early-B and O-type companions (in other words, caused by the Malmquist bias; Malmquist 1922).
Regarding the distribution of the stars excluded from our sample (see Sect. 2.1), we expect Be stars to only populate the lower panels of Fig. 3, since classic Be stars mostly comprise early-B MS objects with luminosity classes III to V, and to do it evenly with (see, e.g., Arcos et al. 2018). On the other hand, SB2+ systems represent an important fraction in the O-star domain. For example, Holgado et al. (2022) showed that 30% of their sample of O-type stars are SB2 systems (see also Barbá et al. 2017; Sana 2017), while in de Burgos et al. (2023, see Fig. 6) we found that the percentage among early- to mid-B-type supergiants drops below 10%. The lack of SB2+ systems could for instance explain the apparently lower number of objects above 30 kK shown in Fig. 2 (see also Sect. 5.2). In addition, the lower number of objects around 32 kK with ¿ 5.0 dex (more clearly shown in Fig. 3) could be explained considering that O9-type stars, which populate that parameter space, gather the largest number of detected SB2 systems. However, a proper diagnostic of this feature would also require a better understanding on how binary interaction products populate the HR diagram. We also note that the left-skewed distributions shown in Fig. 3, particularly for stars with ¡ 5.0 dex, partially result from the effect of stars reaching different maximum values along the ZAMS.